OpenStudy (anonymous):

A Country's population in 1993 was 107 million. In 1999 it was 112 million. Estimate the population in 2017 using the exponential growth formula. Round your answer to the nearest million.

jhonyy9 (jhonyy9):

from 1993 till 1999 is 6 years so from 107 million till 112 million is 5 million so this mean that in 6 years the population numbers growth by 5 million from 1999 till 2017 is 18 years so than in 6 years was the growth 5 million so than in 18 years will be x million ------------------------------------------ x = 18*5/6 = 3*5 = 15 so this mean that the population numbers in 18 years will growth by 15 million than result that in 2017 will be 112 +15 = 127 so will be 127 million